# Aryaman Arora  Neil Rathi  Nikil Roashan Selvam  Róbert Csordás Dan JurafskyChristopher Potts

### v2

@@ 3 Synthetic retrieval tasks
3.1 Associative Treecall
Since a standard AR document (section 2) consists of adjacent key–value pairs, one can associate each key with its corresponding value solely using relative position. Yet many natural language retrieval tasks require association over latent hierarchical structure. For example:
@@ 3.1 Associative Treecall
This type of retrieval is widely studied in cognitive science as binding. The mechanisms underlying natural-language binding in neural networks/LLMs have been examined by Greff et al. [2020], Kim and Schuster [2023], Feng and Steinhardt [2024], Prakash et al. [2024], Li et al. [2025a], Prakash et al. [2025]. Yet no synthetic analogue of this task exists to isolate this mechanism and enable direct comparison to AR. ATR thus allows us to study how different architectures implement binding, and ask if these solutions generalize from simple AR.
@@
An ATR corpus is drawn from a synthetic probabilistic context-free grammar (PCFG) whose parameters we set. Each document consists of a string sampled from the PCFG, with latent structure made up of parent–child relations between symbols, followed by a divider token (EOS) and a query about one such relation. The PCFG has one special property which establishes the parent–child relationships: for the right-hand side of each production rule, the rightmost symbol is always a terminal, and is the parent of the symbols created by this production. We sample strings by selecting an iid nonterminal and recursively applying production rules according to the PCFG distribution. We show an example in Figure 1 and formalise definitions in appendix A. Since the number of tokens separating parents and their children may vary, ATR cannot be solved by a positional associative mechanism.
3.2 Parameters
Param. Description
$H$ Is the head terminal at the left or the right of each production?
$d_{\textrm{max}}$ Maximum depth permitted for the PCFG to generate.
Table 1: Parameters used for constructing a PCFG. We define PCFGs in Greibach Normal Form (GNF); see Appendix A for more details.
PCFG setup.
@@ 3.2 Parameters
For each experiment, we generate a single PCFG to use across all models to ensure fair comparisons, with parameters in Table 1. We also reject any samples that have more than $1024$ symbols, which only affects the sampling distribution for the most complex PCFGs we use.
Queries.
@@
Each PCFG sample of length $n$ provides us with a set of $n-1$ eligible parent–child queries (i.e. a tree with $n-1$ edges). However, terminals may occur multiple times, so a query about a specific symbol may present ambiguity; thus, when presenting a query we consider it to only refer to the rightmost instance of that symbol. [^2] This is the same setup as AR with rewrites Rodkin et al. [2025]. Therefore, the maximum number of eligible queries over all samples is $\min(n-1,\lvert\Sigma\rvert)$. To minimise the ability to heuristically guess, we inversely weight parent–child pairs by the parent’s child count when sampling queries.
3.3 Methodology
Datasets.
We generate synthetic pretraining and evaluation datasets for both tasks. For each setting, the trainset has $100,032$ examples and the eval/dev sets have $320$ examples. In AR, we use disjoint key and value vocabularies; in ATR, keys and values are both sampled from the set of terminals. In each document, we separate the document from the query with a divider token, and provide only a single query. Example AR/ATR documents are in Figure 1; further details in appendix C.
Models.
We pretrain models from scratch on a variety of synthetic tasks. We use the exact architecture implementations from the zoology [^3] https://github.com/HazyResearch/zoology library [Arora et al., 2024b], except for behaviour-preserving modification of the LM backbone to enable interventions with pyvene [^4] https://github.com/stanfordnlp/pyvene [Wu et al., 2024] on the sequence mixers, MLPs, and layer blocks. The LM backbone for all architectures is the same, with pre-norm blocks of alternating sequence mixers and MLPs (except for Mamba, which has no MLP) followed by LayerNorm at the end. We experiment with the following architectures: Attention [Vaswani et al., 2017], BaseConv Arora et al. [2024a], Based Arora et al. [2024b], H3 Fu et al. [2023], Hyena Poli et al. [2023], and Mamba Gu and Dao [2024]; further details on model configurations are given in appendix B.
Training.
We minimise cross-entropy loss, and mask the loss on all tokens except the query (the underlined token in the example below). We use the AdamW optimiser with $\beta=(0.9,0.999),\epsilon=10^{-8}$ and no weight decay. We warm up learning rate for the first $10\%$ of training and then follow a cosine decay schedule to 00 for the remainder of training. We train for either $16$ epochs (on AR) or $32$ epochs (on ATR) with a batch size of $32$. Each experiment trains $\approx 200$ models over all hyperparameters. Runtime varies from $0.5$ to $5$ hours, depending on hardware, task, and architecture. Overall, we used $<10,000$ GPU-hours in total, on a cluster with various NVIDIA machines (with GPU memory ranging from $12.3$G to $143.8$G).
Behavioural metrics.
We report behavioural metrics given the model’s predicted probabilities over the vocabulary $\hat{\mathbf{y}}\in\mathbb{R}^{\lvert\Sigma\rvert}$ and the index of the single true answer $i$. Our main metric is accuracy: $\mathds{1}[\mathrm{arg\,max}(\hat{\mathbf{y}})=i]$. Additionally, we compute but do not primarily report likelihood $\hat{\mathbf{y}}_{i}$.
4 Mechanistic metrics for AR and ATR
Figure 2: Our interchange intervention setup for analysing AR and ATR. Left: We intervene on input and output representations of whole blocks (1 and 6), sequence mixers (e.g. attention blocks; 2 and 3), and state mixers (4 and 5). Right: An example intervention on AR where we corrupt and attempt to restore the key (A) by intervening at the value token in an intermediate representation. We evaluate the downstream effect on the next-token prediction at the query.
Behavioural metrics provide little information as to why certain architectures succeed or fail on tasks of interest. Mechanistic metrics, which directly measure how information flows across model components and token positions, can tell us how AR and ATR are being solved by different architectures, and thus help us understand failures. We illustrate our approach in Figure 2.
We use interchange interventions [Geiger et al., 2021, 2024] to understand and measure how solutions to AR and ATR are implemented across architectures. We introduce this operation and define the resulting metrics for our tasks below. Our implementation uses the pyvene library Wu et al. [2024].
Interchange intervention.
Consider a language model $p(\cdot)$ and some input $\mathbf{b}$. We select a component $f$ inside that model which computes some internal representation $f(\mathbf{b})$ during the LM’s forward pass. Now, consider a counterfactual input $\mathbf{s}$: this produces a counterfactual representation $f(\mathbf{s})$ when processed by $f$. We want to understand what about the output of $p$ is dependent on $f$. Therefore, we perform an intervention which replaces the output $f(\mathbf{b})$ with that of $f(\mathbf{s})$ during the computation of $p(\mathbf{b})$, with the change propagating downstream. The result is notated $p_{f\leftarrow f^{*}}(\mathbf{b},\mathbf{s})$.
Concrete setup for AR and ATR.
We take $\mathbf{o}$ to be a ground-truth document from our data distribution and $\mathbf{c}$ to be a version of that document with exactly one important token corrupted: the key (see Figure 2). This corruption significantly reduces task accuracy for both AR and ATR by removing information that is necessary to answer the query.
We intervene at both the input and output each of the following model components $f$: each layer block, each sequence-mixer, and each state-mixer (i.e. MLP, except in Mamba which lacks this component); see Figure 2, left. We measure to what extent the intervention can restore the likelihood of the correct answer to the query, i.e. we compare restored likelihood $p_{f\leftarrow f^{*}}(y_{\textrm{true}}\mid\mathbf{c},\mathbf{o})$ with original likelihood $p(y_{\textrm{true}}\mid\mathbf{o})$ and corrupted likelihood $p(y_{\textrm{true}}\mid\mathbf{c})$.
Metrics.
Given the above three quantities, we compute attribution score, or what proportion of the original likelihood was restored by the intervention:
$\mathrm{Attrib}(f)$ $=\frac{p_{f\leftarrow f^{*}}(y_{\textrm{true}}\mid\mathbf{b},\mathbf{s})-p(y_{\textrm{true}}\mid\mathbf{b})}{p(y_{\textrm{true}}\mid\mathbf{s})-p(y_{\textrm{true}}\mid\mathbf{b})}$ (1)
For AR and ATR in particular, there are two choices for $f$ which help us distinguish the mechanism underlying task success. To check whether induction is the underlying mechanism, we compute metrics for $f$ being the layer $1$ block input at the value token. Alternatively, we check whether other tokens at layer $1$ block input mediate information flow, indicating some sort of association-less direct retrieval mechanism: the key, query, and divider.
5 Experiments
We now deploy our mechanistic metrics (section 4) on both AR and ATR (section 3). We follow the methodology outlined in section 3.3 to create a variety of AR and ATR datasets and train models with various architectures and hyperparameter configurations. See appendix D for additional experiments not included here.
5.1 (Most) SSMs do not learn induction to solve AR
(a) Model dimensionality vs. accuracy after tuning learning rate for each setting.
(b) Attribution score when restoring the key at its value (induction) vs. at the query, for all LRs.
(c) Same as left but with restoring at the query (layer 0 direct retrieval) vs. key (layer 1).
Figure 3: Associative recall: Accuracy and interchange intervention results on AR with vocabulary size $8192$ and key–value count of $32$. SSMs (except for Based) and Transformers learn different mechanisms.
We run experiments on a relatively simple AR task and show that interchange interventions empirically confirm the same mechanisms underlying AR as proposed in existing theoretical work. We fix the total number of unique keys and values in the vocabulary to be $8192$, and present $32$ key–value pairs in context. Our trainset includes $100032$ examples. We vary model dimensionality in $\{16,32,64,128,256\}$ and sweep LR in the range $[3\cdot 10^{-5},3\cdot 10^{-2}]$ for each architecture.
@@ 5 Experiments
Figure 3(a) demonstrates that task accuracy on AR cleanly separates Attention, which achieves $100\%$ accuracy at $d\geq 32$, from nearly all SSMs. Based solves AR near-perfectly with roughly the same dimension-wise scaling curve as Attention, achieving a maximum accuracy of $99.06\%$. However, Mamba is a close third and clearly better than other SSMs at AR, albeit achieving a less-than-perfect $91.25\%$ at $d=256$.
@@
[…]
SSMs perform direct retrieval at varying layers: the best-performing Mamba, Hyena, and H3 models almost entirely perform direct retrieval at layer 0 via the query token, while worse SSM checkpoints use a mix of query and key tokens, indicating delayed direct retrieval by both layer 0 and layer 1. Jelassi et al. [2024] shows that direct retrieval in SSMs has asymptotically worse capacity than the induction solution, and this is reflected in performance on AR.
5.2 Per-architecture mechanisms are similar between ATR and AR
We consider four initial settings to study models on ATR, over all combinations of $L_{\textrm{max}}=\{5,10\}$ and $\lvert\Sigma\rvert=\{20,8192\}$. We keep all other parameters fixed with settings given in appendix C. Varying $L_{\textrm{max}}$ controls the possible distances between keys and values in the PCFG sample without affecting other properties that play a role in task difficulty (e.g. depth). Varying $\lvert\Sigma\rvert$ stresses the state capacity, since more key–value pairs must be tracked, without affecting syntactic complexity. We sweep the same model dimensionalities as in section 5.1, and a smaller learning rate range of $[3\cdot 10^{-5},3\cdot 10^{-3}]$.
Behavioural results.
We report results in Figure 4(a). Surprisingly, Mamba is highly successful at ATR. On the small terminal count setting ($\lvert\Sigma\rvert=20$) Mamba matches or outperforms all other architectures at all model dimensions, particularly with longer production rules ($L_{\textrm{max}}=10$) with performance of $92.19\%$ vs. $80.94\%$ for Attention at $d=256$. This is particularly surprising because longer production rules imply greater positional variation between keys and values, which ought to stress AR-focused SSM designs. Attention only manages to outperform Mamba slightly on the large terminal count setting ($\lvert\Sigma\rvert=8192$) when $d\leq 64$.
Mechanistic analysis.
@@ 5.2 Per-architecture mechanisms are similar between ATR and AR
We conduct the same analysis as for AR. We recover the same overall trends but with greater inter-architecture variance: Figure 4(b) shows that Attention, Based, and BaseConv all primarily learn induction mechanisms, whereas the remaining SSMs perform direct retrieval as on AR, with high attribution scores on either the key (indicating direct retrieval by the layer 1 sequence mixer) or the query (indicating the same but by layer 0).
Intriguingly, Figure 4(b) shows that different SSMs form different strategies across task difficulties; in particular, all direct-retrieval SSMs favour delaying retrieval to layer 1 when terminal count is large ($\lvert\Sigma\rvert=8192$), but use a mix of layers otherwise. Regardless, the same tendency from AR recurs: SSMs besides Based and BaseConv do not perform induction, but Mamba is still highly performant. Strikingly, as the next section shows, Mamba also achieves high generalization performance on ATR.
(a) Model dimensionality vs. accuracy after tuning learning rate for each setting.
(b) Summarised attribution scores at key, value, and query for each setting, when restoring the key.
Figure 4: Associative Treecall: Accuracy and interchange intervention results on ATR across varying settings. The same trend as on AR holds, with Attention, Based, and Mamba achieving high performance but with entirely different mechanisms.
5.3 Mamba’s solution to ATR does generalise
We reuse the easiest settings from our ATR experiment ($L=5,\lvert\Sigma\rvert=20$) and construct a new dataset with a train–test split on query–answer pairs. Specifically, $80\%$ of possible unique query–answer pairs are provided in the training set, while $20\%$ are only in the test set and thus never trained on. We seek to assess whether models learn a general mechanism for parent–child relations in ATR or if the impressive results of Mamba (as well as Attention and Based) are merely the result of better memorisation of the PCFG parameters. This setup is akin to Wang et al. [2024]’s technique of train–test split on multi-hop queries; we provide supervision on individual query and answer types, but not on some compositions of them.
Behavioural results.
We select the checkpoint with the highest dev accuracy for each architectural and dimensionality setting, after sweeping LR. We plot the dev and test accuracies of each of these checkpoints in Figure 5(a); all models have much lower test accuracy (e.g. Attention with $d=256$ has $95.62\%$ dev and $68.12\%$ test accuracy). Attention achieves the greatest dev accuracies on $d\geq 32$. Mamba’s relative ranking is lower than on the in-distribution setting in section 5.2, but it still achieves the overall second-highest dev accuracy ($65.00\%$ at $d=128$). Surprisingly, H3 generalises well despite its poor dev accuracy, beating Mamba on test accuracy in $3$ out of $5$ settings.
We compare dev and test accuracies across all LRs in Figure 5(b). We find that while Mamba does have unusually high dev accuracy given a selected test accuracy (indicating greater memorisation than models with other architectures), its dev accuracy is still generally higher than non-Attention architectures. Interestingly, H3 has nearly Attention-level generalisation while BaseConv exhibits vanishingly little generalisation. Overall, behavioural metrics show that Mamba does nontrivially generalise on ATR, albeit not as well as Attention.
Mechanistic analysis.
We report a summary of attribution scores at different tokens (key, query, value), comparing on dev and test sets across all checkpoints in Figure 5(c). We find largely consistent mechanisms underlying behaviour on both dev and test, and these match attribution scores on ATR without train–test split. The only exception is that BasedConv does induction on the dev set but not nearly as much on the test set; its induction mechanism is more brittle than Attention and Based.
5.4 Short convolutions enable AR and ATR in Mamba and Based
Throughout all our experiments on AR and ATR, we repeatedly observed that Attention, Based, and Mamba are the highest-performing architectures. However, their underlying mechanisms differ: Attention and Based learn induction, a 2-layer mechanism which stores key–value associations at the value token as an intermediate step, whereas Mamba uses direct retrieval, a 1-layer mechanism which directly writes an association to the query token.
Importantly, Based and Mamba share a key architectural component: short convolutions. We hypothesise that this component is necessary [^6] Since Hyena also has a short convolution, this may not be sufficient for good performance on association. for performing association (as in AR and ATR) when using a subquadratic sequence mixer. We conduct experiments on AR where we shorten the convolution kernel size in Mamba (from the default $d_{\mathrm{conv}}=4$ to $\{3,2,1\}$, and deleting it) and replace the Based short convolution with implicitly-parametrised long convolution [Poli et al., 2023].
@@ Results.
We report results of our ablations in Figure 6. On Mamba (Figure 6(a)), we find a step change in task accuracy when increasing $d_{\mathrm{conv}}$ from $1$ to $2$, which introduces previous token information and thus enables AR. Without short convolution, Mamba fails to learn AR. Figure 6(b) further shows that larger kernel size leads to earlier (in layer 0) direct retrieval. Finally, besides $d_{\mathrm{conv}}<2$ like Mamba, implicit long convolution in Based also significantly harms AR performance (Figure 6(c)). Therefore, we conclude that short convolutions are responsible for association on AR in Mamba and Based.
(a) Accuracy on AR (length $32$) for Mamba when varying the kernel size of the short convolution.
(b) Summarised attribution scores across all checkpoints for Mamba when varying conv. kernel size.
(c) Accuracy on AR for Based when varying conv. kernel size or using implicit long conv.
Figure 6: Ablating short convolution: Accuracy and interchange intervention results when ablating parameters of the short convolution component in Mamba, Based, and BaseConv.
@@ 6 Discussion
Architectural advances on language modelling are largely uncovered and presented in an empirical manner; beyond intuition, we have little justification as to why a modification or innovation improves model performance. Synthetic tasks already regularly inform progress on subquadratic architecture design (such as SSMs), but treating such tasks as another downstream evaluation is loses useful signal; control over task parameters presents an opportunity to explain performance using interpretability.
@@
We introduced ATR to break the naïve key–value adjacency of AR, and see whether general mechanisms underlying association still emerge across architectures. We find the same induction mechanism, where the association is computed and stored at the value before retrieval, in Attention and Based for both tasks. While Olsson et al. [2022] and later works define induction on adjacent tokens, ATR is evidence that a position-independent and generalising (section 5.3) notion of association can be implemented by a single attention head. Further investigation of ATR (e.g. multi-hop queries) is necessary to understand the limits of induction.
@@
We showed that Mamba and Based rely on short convolutions to learn how to associate keys and values on AR and ATR. Several earlier works point to the importance of short convolution: Arora et al. [2024b] empirically show its utility on AR (along with sliding-window attention), Allen-Zhu and Alfarano [2025] introduce a short convolution component (Canon) in various architectures to improve synthetic and real task performance, and Olsson et al. [2022] show that 1-layer attention can learn induction if augmented with a length-2 convolution; further see Liu et al. [2024b], Dolga et al. [2024], Fu et al. [2023], Poli et al. [2023].
7 Limitations
While we proposed mechanistic evaluations as a new tool, behavioural metrics like accuracy are still needed to properly contextualise results. Additionally, here we did not perform mechanistic evaluation of subcomponents of sequence mixers (e.g. the selective SSM component within Mamba), due to implementation difficulties when applying interventions within hardware-optimised operators, which are inaccessible via PyTorch hooks. Finally, we focus on synthetic tasks throughout this work; extending our analyses to real-world models would help paint a more complete picture of the differences in capabilities (and underlying mechanisms) of different architectures on real-world tasks.

### v3

@@ 5.1 Associative Treecall (ATR)
Since a standard AR document (section˜2) consists of adjacent keyvalue pairs, one can associate each key with its corresponding value solely using relative position. Yet many natural language retrieval tasks require association over latent hierarchical structure. For example:
@@
An ATR corpus is drawn from a synthetic probabilistic context-free grammar (PCFG) whose parameters we set. Each document consists of a string sampled from the PCFG, with latent structure made up of parent–child relations between symbols, followed by a divider token (EOS) and a query about one such relation. The PCFG has one special property which establishes the parent–child relationships: for the right-hand side of each production rule, the rightmost symbol is always a terminal, and is the parent of the symbols created by this production. We sample strings by selecting an iid nonterminal and recursively applying production rules according to the PCFG distribution. We show an example in Figure˜6 and formalise definitions in appendix˜B. Since the number of tokens separating parents and their children may vary, ATR cannot be solved by a positional associative mechanism.
Setup.
@@
For each experiment, we generate a single PCFG to use across all models to ensure fair comparisons, with parameters in Table 2. We also reject any samples that have more than $1024$ symbols, which only affects the sampling distribution for the most complex PCFGs we use. We follow section˜4.1 except we train for $32$ epochs.
@@
Each PCFG sample of length $n$ provides us with a set of n1n-1 eligible parent–child queries (i.e. a tree with n1n-1 edges). However, terminals may occur multiple times, so a query about a specific symbol may present ambiguity; thus, when presenting a query we consider it to only refer to the rightmost instance of that symbol. To minimise the ability to heuristically guess, we inversely weight parent–child pairs by the parent’s child count when sampling queries.
5.2 Per-architecture mechanisms are similar between ATR and AR
We consider a simple setting to study models on ATR: the maximum number of symbols on the right-hand side of any rule $L_{\textrm{max}}=5$ and vocabulary size $\lvert\Sigma\rvert=20$. The remaining parameters are set in accordance with appendix˜C; see Table˜2 for definitions. We sweep the same model dimensionalities as in section˜4.2, and a smaller learning rate range of $3\cdot 10^{-5},3\cdot 10^{-3}]$.
@@ 5.2 Per-architecture mechanisms are similar between ATR and AR
We report results in the leftmost panel of Figure˜7. Mamba is the best-performing architecture at ATR; it matches or outperforms all other architectures at all model dimensions. This is particularly surprising because longer production rules imply greater positional variation between keys and values, which ought to stress AR-optimised SSM designs like Mamba.
Unlike AR, the baseline performance on ATR after corrupting the key is well above 0. This is because some keyvalue pairings are more likely than others due to the underlying PCFG, enabling memorisation (unlike AR where all pairings are equally likely).
@@
We conduct the same analysis as for AR. We recover the same overall trends but with greater inter-architecture variance: Figure˜7 shows that Attention, Based, and BaseConv all primarily learn induction mechanisms with performance largely restored by value interventions, whereas the remaining SSMs perform direct retrieval as on AR, with performance being restored by intervening on the key (indicating direct retrieval by the layer 1 sequence mixer) or the query (indicating the same but by layer 0).
Figure 7: Associative Treecall: Likelihood of correct answer without any interventions (leftmost), after corrupting the key (second), and after restoring representations at the layer 1 block input (at key, value, or query) with interchange intervention. The pattern from AR largely holds.
5.3 Mamba can learn ATR without short convolutions
Figure 8: Ablating short convolution in Mamba on ATR: Original, corrupted key, and restored interventions at layer 1 input at the query, key, and value in Mamba, when varying model dimensionality and conv. size. (Some checkpoints failed to train.)
Given ATR and AR show similar patterns in mechanistic implementation across architectures, we now ask whether Mamba’s reliance on short convolution holds on this new task. Since ATR may have longer and more variable distances between the key and value than AR, short convolution (with a short receptive field that applies a fixed input-unaware transformation) may not be the appropriate operator for performing association.
Methodology.
We repeat the experiments in section˜4.3 on ATR, training Mamba with varying conv. size or without conv. We apply mechanistic metrics to the best-performing checkpoints of each dimensionality and conv. setting.
@@ Results.
We report results in Figure˜8. Unlike in AR, we find that Mamba still learns the task even without short convolutions (but not as well as with them). E.g. at $d=256$, Mamba without short convolutions ($93.4\%$) still outperforms Based ($92.2\%$) and BaseConv ($89.7\%$). The more surprising result is via mechanistic metrics: without short convolutions or with $d_{\text{conv}}=1$, Mamba’s performance after corruption can be restored primarily by intervening at the value, i.e. Mamba without short convolutions learns two-layer induction on ATR. We have thus shown that Mamba can indeed learn an induction-like mechanism to solve associative recall tasks, albeit using different components than Attention and only on certain tasks (ATR, but not AR). This may explain other induction-like mechanisms observed in Mamba trained on natural language, e.g. Bick et al. (2025), and highlights a divergence between AR and real-world retrieval.
@@ 6 Discussion
Architectural advances on language modelling are largely uncovered and presented in an empirical manner (e.g. Shazeer, 2020); beyond intuition, we have little justification as to why a modification or innovation improves model performance. Synthetic tasks already inform progress on architecture design (such as SSMs), but treating such tasks as just a behavioural evaluation discards useful signal; control over task parameters presents an opportunity to explain performance using interpretability.
@@
We introduced ATR to break the naïve key–value adjacency of AR, and see whether general mechanisms underlying association still emerge across architectures. We find the same induction mechanism, where the association is computed and stored at the value before retrieval, in Attention and Based for both tasks, as well as in Mamba without convs. While Olsson et al. (2022) and later works define induction on adjacent tokens, ATR is evidence that a position-independent and generalising (section˜D.2) notion of association is worth studying.
@@
We showed that Mamba and Based rely on short convolutions to learn how to associate keys and values on AR and ATR. Several earlier works point to the importance of short convolution: Arora et al. (2024b) empirically show its utility on AR, Allen-Zhu and Alfarano (2025) introduce a short convolution component (Canon) in various architectures,Olsson et al. (2022) show that 1-layer attention can learn induction if augmented with a length-2 convolution; also Liu et al. (2024b); Dolga et al. (2024); Fu et al. (2023); Poli et al. (2023).
Limitations.
While we proposed mechanistic evaluations as a new tool, behavioural metrics like accuracy are still needed to properly contextualise results. Additionally, we focus on synthetic tasks throughout this work; extending our analyses to real-world models would help paint a more complete picture of the differences in capabilities and mechanisms of architectures on real-world tasks.